# Fusions of One-Variable First-Order Modal Logics

### 04 Mar 2026 v1

@@ 4 Fusions of One-Variable Logics with Equality
be the (non-terminating) computation of $M$ on $(0,0)$. We define a (cactus-like) model $\mathfrak{M}$ over a frame $\mathfrak{F}\in\mathscr{C}_{1}\otimes\mathscr{C}_{2}$ with a constant domain $\mathbb{N}$ such that $\mathfrak{M}\models\varphi_{Q}\wedge\varphi_{I}$, but $\mathfrak{M},w_{0}\models\varphi_{\mathit{init}}$, for some world $w_{0}$.
Let $\mathfrak{G}_{1}=(U_{1},S_{1})$ and $\mathfrak{G}_{2}=(U_{2},S_{2})$ be non-trivial frames from, respectively, $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$; thus, $\mathfrak{G}_{1}$ contains worlds $w$ and $v$ such that $w\neq v$ and $wS_{1}v$; similarly, $\mathfrak{G}_{2}$ contains worlds $s$ and $t$ such $s\neq t$ and $sS_{2}t$. The frame $\mathfrak{F}=(W,R_{1},R_{2})\in\mathscr{C}_{1}\otimes\mathscr{C}_{2}$ we use is a limit cactus constructed from a sequence $\{\mathfrak{G}_{1}^{i}=(U_{1}^{i},S_{1}^{i})\mid i<\omega\}$ of disjoint copies of $\mathfrak{G}$ and a sequence $\{\mathfrak{G}_{2}^{i}=(U_{2}^{i},S_{2}^{i})\mid i<\omega\}$ of disjoint copies of $\mathfrak{G}_{2}$ (starting with $\mathfrak{G}^{0}_{1}$). If $u\in U_{k}$, for $k=1,2$, then its copy from $U^{i}_{k}$ is denoted by $u^{i}$.
To simulate the computation $C_{0}$, we shall use the infinite sequence
$w^{0}R_{1}v^{0}(=s^{0})R_{2}t^{0}(=w^{1})R_{1}v^{1}(=s^{1})R_{2}\ldots$ (1)
of worlds of $\mathfrak{F}$. The interpretation of non-logical symbols at $w^{k}$ will simulate the $k$th configuration of $C_{0}$. Since $v$ might not be a unique successor of $w$ in $\mathfrak{G}_{1}$, and $t$ might not be a unique successor of $s$ in $\mathfrak{G}_{2}$, we will make sure that interpretation of non-logical symbols over each sequence of words of the form $w^{0}R_{1}u_{1}R_{2}u_{2}R_{1}u_{3}R_{2}\ldots$ agrees with the interpretation of those symbols over the sequence (1).
We define, by recursion, an infinite sequence $W^{0},W^{1},\ldots$ of subsets of $W$ and the interpretation of non-logical symbols at worlds from those sets in such a way that the interpretation of all non-logical symbols at all worlds from $W^{n}$, for a fixed $n$, is the same.
First, define $W^{0}=\{w^{0}\}$ and let $I_{w^{0}}(V_{1})=I_{w^{0}}(V_{2})=\emptyset$; $I_{w_{0}}(q_{0})=I_{w_{0}}(p)=1$; $I_{w_{0}}(q_{i})=0$, for all $i\geq 1$; $I_{w_{0}}(c)\in\mathbb{N}$ be arbitrary. Then, $w^{0}\models\varphi_{\mathit{init}}$.
Suppose that the set $W^{n}$, with $n\geq 0$, and the interpretation of non-logical symbols at worlds of $W^{n}$ have been defined so that this interpreation is the same at every world from $W^{n}$. Then, put
$\begin{array}[]{lcl}W^{n+1}&=&\{w\in W\mid\exists u\in W^{n}\,\exists v\in W\,(uR_{1}vR_{2}w~\&~u\neq v~\&~v\neq w)\}.\end{array}$
For every $w\in W^{n+1}$, put (below, $v$ is an arbitrary world from $W^{n}$; the choice does not matter due to IH):
@@
  • $I_{w}(q_{j})=1$ iff $q^{n+1}=q_{j}$;
  • $I_{w}(V_{1})$ and $I_{w}(V_{2})$ depend on the instruction $I^{n}$ executed at step $n$ of $C_{0}$:
    • if $I^{n}=+(r_{k},q_{j})$, then $I_{w}(V_{k})=I_{v}(V_{k})\cup\{I_{v}(c)\}$ and $I_{w}(V_{3-k})=I_{v}(V_{3-k})$;
    • if $I^{n}=-(r_{k},q_{j},q_{l})$ and $v_{k}\neq 0$, then $I_{w}(V_{k})=I_{v}(V_{k})\setminus\{I_{v}(c)\}$ and $I_{w}(V_{3-k})=I_{v}(V_{3-k})$;
    • if $I^{n}=-(r_{k},q_{l},q_{j})$ and $v_{k}=0$, then $I_{w}(V_{k})=I_{v}(V_{k})$ and $I_{w}(V_{3-k})=I_{v}(V_{3-k})$;
  • $I_{w}(c)$ depends on the instruction $I^{n+1}$ executed at step $n+1$ of $C_{0}$:
    • if $I^{n+1}=+(r_{k},q_{j})$ or $I^{n+1}=-(r_{k},q_{j},q_{l})$ and $v_{k}=0$, then $I_{w}(c)\notin I_{w}(V_{k})$;
    • if $I^{n+1}=-(r_{k},q_{l},q_{j})$ and $v_{k}\neq 0$, then $I_{w}(c)\in I_{w}(V_{k})$.
This completes the definition of the $W^{i}$ and the values of $I$ at worlds from $\bigcup\nolimits_{i<\omega}W^{i}$. If $w\in W\setminus\bigcup\nolimits_{i<\omega}W^{i}$, then we put $I_{w}(p)=0$ and $I_{w}(q_{i})=0$, for every $i\in\{1,\ldots,N\}$; the interpretation of other non-logical symbols at such worlds is immaterial. Let $\mathfrak{M}=(\mathfrak{F},\langle\mathbb{N},I_{w}\rangle_{w\in W})$. Then, $\mathfrak{M}\models\varphi_{Q}\wedge\varphi_{I}$.
$(\Leftarrow)$ Suppose there is a model $\mathfrak{M}=((W,R_{1},R_{2}),\langle\Delta,I_{w}\rangle_{w\in W})$ such that $\mathfrak{M}\models\varphi_{Q}\wedge\varphi_{I}$ but $\mathfrak{M},w_{0}\models\varphi_{\mathit{init}}$, for some $w_{0}\in W$. We define an infinite sequence $(q^{0},v^{0}_{1},v^{0}_{2}),\,(q^{1},v^{1}_{1},v^{1}_{2}),\,(q^{2},v^{2}_{1},v^{2}_{2}),\,\ldots$ of configurations of a Minsky machine and show that this sequence is the computation of $M$ on $(0,0)$.
First, let
  • $q^{0}=q_{i}$ iff $\mathfrak{M},w_{0}\models q_{i}$;
  • $v^{0}_{k}=|I_{w_{0}}(V_{k})|$, for $i=1,2$.
Since $\mathfrak{M},w_{0}\models\varphi_{\mathit{init}}$ and $\mathfrak{M}\models\varphi_{Q}$, it follows that $(q^{0},v^{0}_{1},v^{0}_{2})$ is the initial configuration on $(0,0)$. From there, we can follow formulas $\varphi_{i}$ to construct the non-terminating computation of $M$ on $(0,0)$. ∎
@@ 5 Fusions of Propositional Modal Logics Sharing an S5 Modality
Let $L_{1}$ and $L_{2}$ admit $E$-homogeneous models (for global consequence). Then $L_{1}\otimes_{E}L_{2}$ is
  • (globally) Kripke complete;
  • (globally) decidable if $L_{1}$ and $L_{2}$ are both (globally) decidable.
Modulo replacing $\forall x$ with $\Box_{E}$, we use the same notation as in Section 3. For instance, if $\Phi$ is a finite set of $\mathcal{L}_{1,2,E}$-formulas closed under subformulas, then a $\Phi$-type is a maximal Boolean-consistent subset of $\{\psi,\neg\psi\mid\psi\in\Phi\}$, and a $\Phi$-quasistate is a non-empty set of $\Phi$-types satisfying the $\Box_{E}$-counterpart of (t). We set
@@
We now inductively define a tapered $(i,\kappa)$-cactus $\mathfrak{C}$ for $\Phi$ as a tuple of the form $((W,R_{1},R_{2},E),T_{0},\tau)$, where $(W,R_{i},E)\models L_{i}$, $T_{0}\subseteq W$ is an $E$-equivalence class, no node in $T_{0}$ is in the domain or range of $R_{3-i}$, $(W\setminus T_{0},R_{2},E\setminus T_{0}\times T_{0})\models L_{3-i}$, and $\tau$ is a tapered mapping such that, for every set $\Xi$ of formulas and $E$-equivalence class $T$, $|\{w\in T\mid\tau(w)=\Xi\}|\in\{0,\kappa\}$ and which is $1$-, 22- and $E$-coherent, and $i$- and $E$-saturated on $W$ but (3i)(3-i)-saturated only on $W\setminus T_{0}$.
For the base case, one can easily show that a tapered $(i,\kappa)$-cactus $((W,R_{1},R_{2},E),T_{0},\tau)$ for $\emptyset$ always exists. Note that, as $\Phi=\emptyset$, we have $\tau(w)=\emptyset$ for all $w\in W$ and $|T_{0}|=\kappa$.
@@
A fully grafted tapered $(i,\kappa)$-cactus $((W,R_{1},R_{2},E),\tau)$ for $\Phi$ results from grafting a compatible tapered $((3-i),\kappa)$-cactus $\mathfrak{C}^{T^{0}}$ for $\Phi^{T^{0}}=\Theta^{\mathbf{i}}(\Phi)$ onto $T^{0}$ in a tapered $(i,\kappa)$-cactus $(W^{0},R_{1}^{0},R_{2}^{0},E^{0},T^{0},\tau^{0})$ for $\Phi$. Observe that the resulting mapping $\tau$ is $1$-, 22- and $E$-coherent and $1$-, 22- and $E$-saturated on $W$.
@@
For each $E^{0}$-equivalence class $T$ with $h(T)\geq 0$, consider the $\Theta^{1}(\Phi)$-quasistate $\boldsymbol{q}^{T}\in\boldsymbol{Q}_{1}(\varphi)$ such that $\mathfrak{M}^{0},v\models(\hat{\boldsymbol{q}}^{T})^{\mathbf{1}}$ for all (equivalently, some) $v\in T$. Observe that $\boldsymbol{q}^{T}$ may contain formulas $\chi$ with $\textit{md}^{\mathbf{1}}{(\chi)}>h(T)$. Let $\Phi^{T}=\Theta^{\mathbf{1}}(\Phi)\cap\Gamma^{\mathbf{1}}_{h(T)}(\Phi)$ and consider the restriction $\boldsymbol{q}_{0}^{T}=\{t|_{\smash{\Gamma^{\mathbf{1}}_{h(T)}(\Phi)}}\mid t\in\boldsymbol{q}^{T}\}$, which is a $\Phi^{T}$-quasistate. Denote $\smash{\hat{\boldsymbol{q}}^{T}_{0}}$ by $\varphi^{T}$. Since $\textit{adp}^{\mathbf{2}}(\varphi^{T})=k-1$, we apply our induction hypothesis to the $\mathcal{L}_{1,2,E}$-formula $\varphi^{T}$ and obtain a tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}=((W^{T},R_{1}^{T},R_{2}^{T},E^{T}),T,\tau^{T})$ for $\mathop{\textit{sub}}\varphi^{T}$ (equivalently, for $\Phi^{T}$) such that $\tau^{0}(v)|_{\Theta^{\mathbf{1}}(\Phi)}=\tau^{T}(v)$ for all $v\in T$. At each $E^{0}$-equivalence class $T$ with $h(T)<0$, we take the tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}$ for $\emptyset$. Next, for each $T\neq T_{0}$, we graft the tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}$ onto $T$ in the $(1,\kappa)$-quasimodel $\mathfrak{Q}$ to obtain a tapered $(1,\kappa)$-cactus $\mathfrak{C}$. A fully grafted tapered $(1,\kappa)$-cactus is obtained by grafting, in addition, $\mathfrak{C}^{T^{0}}$ onto $T^{0}$ in $\mathfrak{C}$.

### 06 Mar 2026 v2

@@ 4 Fusions of One-Variable Logics with Equality
be the non-terminating computation of $M$ on $(0,0)$. We define a cactus-like model $\mathfrak{M}$ based on a frame $\mathfrak{F}\in\mathscr{C}_{1}\otimes\mathscr{C}_{2}$ with a constant domain $\mathbb{N}$ such that $\mathfrak{M}\models\varphi_{Q}\wedge\varphi_{I}$ and $\mathfrak{M},w^{0}\models\varphi_{\mathit{init}}$, for some world $w^{0}$.
Let $\mathfrak{G}_{1}=(U_{1},S_{1})\in\mathscr{C}_{1}$ and $\mathfrak{G}_{2}=(U_{2},S_{2})\in\mathscr{C}_{2}$ be non-trivial frames; thus, $\mathfrak{G}_{1}$ contains worlds $w$ and $v$ such that $w\neq v$ and $wS_{1}v$; similarly, $\mathfrak{G}_{2}$ contains worlds $s$ and $t$ such $s\neq t$ and $sS_{2}t$. The frame $\mathfrak{F}=(W,R_{1},R_{2})\in\mathscr{C}_{1}\otimes\mathscr{C}_{2}$ is a limit cactus constructed from a sequence $\{\mathfrak{G}_{1}^{i}=(U_{1}^{i},S_{1}^{i})\mid i\in\mathbb{N}\}$ of disjoint copies of $\mathfrak{G}_{1}$ and a sequence $\{\mathfrak{G}_{2}^{i}=(U_{2}^{i},S_{2}^{i})\mid i\in\mathbb{N}\}$ of disjoint copies of $\mathfrak{G}_{2}$ (starting with $\mathfrak{G}^{0}_{1}$). To simulate the computation $C_{0}$, we shall use infinite sequences of worlds from $\mathfrak{F}$ of the form
$\begin{array}[]{c}w^{0}R_{1}v^{0}(=s^{0})R_{2}t^{0}(=w^{1})R_{1}v^{1}(=s^{1})R_{2}\ldots\end{array}$
where $w^{i},v^{i}\in U^{i}_{1}$ and $s^{i},t^{i}\in U^{i}_{2}$, and also $w^{i}\neq v^{i}$ and $s^{i}\neq t^{i}$, for all $i\in\mathbb{N}$ (by the assumption about non-triviality of $\mathfrak{G}_{1}$ and $\mathfrak{G}_{2}$, the set of such sequences is non-empty). The $n$th configuration of $C_{0}$ will be simulated by the interpretation of non-logical symbols at worlds $w^{n}$.
More precisely, we define, by recursion, an infinite sequence $W^{0},W^{1},\ldots$ of subsets of $W$ and the interpretation of non-logical symbols at worlds from those sets in such a way that the interpretation of all non-logical symbols at all worlds from $W^{i}$, for a fixed $i$, is the same.
Define $W^{0}=\{w^{0}\}$ and,
$W^{i+1}=\bigl\{w\in W\mid\exists u\in W^{i}\ \exists v\in W\,(uR_{1}vR_{2}w~\&~u\neq v~\&~v\neq w)\bigr\},\qquad\text{for all }i\in\mathbb{N}.$
Then, define the interpretation at worlds $w$ from each $W^{i}$ inductively as follows:
@@
  • $I_{w}(q_{j})=1$ iff $q^{i}=q_{j}$;
  • $I_{w^{0}}(V_{1})=I_{w^{0}}(V_{2})=\emptyset$;
  • $I_{w}(V_{1})$ and $I_{w}(V_{2})$, for $w\in W^{i}$ with $i>0$, depend on the instruction $I^{i-1}$ executed at step (i1)(i-1) of $C_{0}$ (below, $v$ is an arbitrary world from $W^{i-1}$; the choice does not matter since the interpretation of symbols is uniform in $W^{i-1}$):
    • if $I^{i-1}=+(r_{k},q_{j})$, then $I_{w}(V_{k})=I_{v}(V_{k})\cup\{I_{v}(c)\}$ and $I_{w}(V_{3-k})=I_{v}(V_{3-k})$;
    • if $I^{i-1}=-(r_{k},q_{j},q_{l})$ and $I_{v}(V_{k})\neq\emptyset$, then $I_{w}(V_{k})=I_{v}(V_{k})\setminus\{I_{v}(c)\}$ and $I_{w}(V_{3-k})=I_{v}(V_{3-k})$;
    • if $I^{i-1}=-(r_{k},q_{l},q_{j})$ and $I_{v}(V_{k})=\emptyset$, then $I_{w}(V_{k})=I_{v}(V_{k})$ and $I_{w}(V_{3-k})=I_{v}(V_{3-k})$;
  • $I_{w}(c)$ depends on the instruction $I^{i}$ executed at step $i$ of $C_{0}$:
    • if $I^{i}=+(r_{k},q_{j})$ or $I^{i}=-(r_{k},q_{j},q_{l})$ and $I_{w}(V_{k})=\emptyset$, then $I_{w}(c)$ is any fixed $I_{w}(c)\notin I_{w}(V_{k})$;
    • if $I^{i}=-(r_{k},q_{l},q_{j})$ and $I_{w}(V_{k})\neq\emptyset$, then $I_{w}(c)$ is any fixed $I_{w}(c)\in I_{w}(V_{k})$.
Lastly, for $w\in W\setminus\bigcup\nolimits_{i\in\mathbb{N}}W^{i}$, we put $I_{w}(p)=0$ and $I_{w}(q_{i})=0$, for every $0\leq i\leq N$; the interpretation of other non-logical symbols at such worlds is immaterial. Then, $\mathfrak{M},w^{0}\models\varphi_{\mathit{init}}$ and $\mathfrak{M}\models\varphi_{Q}\wedge\varphi_{I}$.
$(\Leftarrow)$ Suppose there is a model $\mathfrak{M}=((W,R_{1},R_{2}),\langle\Delta,I_{w}\rangle_{w\in W})$ such that $\mathfrak{M}\models\varphi_{Q}\wedge\varphi_{I}$ and $\mathfrak{M},w^{0}\models\varphi_{\mathit{init}}$, for some $w^{0}\in W$. We define an infinite sequence $(q^{0},v^{0}_{1},v^{0}_{2}),\,(q^{1},v^{1}_{1},v^{1}_{2}),\,(q^{2},v^{2}_{1},v^{2}_{2}),\,\ldots$ of configurations of a Minsky machine and show that this sequence is the computation of $M$ on $(0,0)$. First, let
  • $q^{0}$ be a uniquely determined $q_{i}$, for $0\leq i\leq N$, such that $I_{w^{0}}(q_{i})=1$;
  • $v^{0}_{k}=|I_{w^{0}}(V_{k})|$, for $i=1,2$.
Since $\mathfrak{M},w^{0}\models\varphi_{\mathit{init}}$ and $\mathfrak{M}\models\varphi_{Q}$, it follows that $(q^{0},v^{0}_{1},v^{0}_{2})$ is the initial configuration on $(0,0)$. From there, we can follow formulas $\varphi_{i}$ to construct the non-terminating computation of $M$ on $(0,0)$. ∎
@@ 5 Fusions of Propositional Modal Logics Sharing an S5 Modality
Let $L_{1}$ and $L_{2}$ admit $E$-homogeneous models (for global consequence). Then
  • $L_{1}\otimes_{E}L_{2}$ is (globally) Kripke complete;
  • $L_{1}\otimes_{E}L_{2}$ is (globally) decidable if $L_{1}$ and $L_{2}$ are both (globally) decidable.
Modulo replacing $\forall x$ with $\Box_{E}$, we use the same notation as in Section 3. For instance, if $\Phi$ is a finite set of $\mathcal{L}_{1,2,E}$-formulas closed under subformulas, then a $\Phi$-type is a maximal Boolean-consistent subset of $\{\psi,\neg\psi\mid\psi\in\Phi\}$, and a $\Phi$-quasistate is a non-empty set of $\Phi$-types satisfying the $\Box_{E}$-counterpart of (qs). We set
@@
We now inductively define a tapered $(i,\kappa)$-cactus $\mathfrak{C}$ for $\Phi$ as a tuple of the form $((W,R_{1},R_{2},E),T_{0},\tau)$, where $(W,R_{i},E)\models L_{i}$, $T_{0}\subseteq W$ is an $E$-equivalence class, no node in $T_{0}$ is in the domain or range of $R_{3-i}$, $(W\setminus T_{0},R_{3-i},E\setminus T_{0}\times T_{0})\models L_{3-i}$, and $\tau$ is a tapered mapping such that, for every set $\Xi$ of formulas and $E$-equivalence class $T$, $|\{w\in T\mid\tau(w)=\Xi\}|\in\{0,\kappa\}$ and which is $1$-, 22- and $E$-coherent, and $i$- and $E$-saturated on $W$ but (3i)(3-i)-saturated only on $W\setminus T_{0}$.
For the base case, one can easily show that a ‘degenerate’ tapered $(i,\kappa)$-cactus $((W,R_{1},R_{2},E),T_{0},\tau)$ for $\emptyset$ always exists. Note that, as $\Phi=\emptyset$, we have $\tau(w)=\emptyset$ for all $w\in W$ and $|T_{0}|=\kappa$.
@@
A fully grafted tapered $(i,\kappa)$-cactus $((W,R_{1},R_{2},E),\tau)$ for $\Phi$ results from grafting a compatible tapered $((3-i),\kappa)$-cactus $\mathfrak{C}^{T^{0}}$ for $\Phi^{T^{0}}=\Theta^{\mathbf{i}}(\Phi)$ onto $T^{0}$ in a tapered $(i,\kappa)$-cactus $((W^{0},R_{1}^{0},R_{2}^{0},E^{0}),T^{0},\tau^{0})$ for $\Phi$. Observe that the resulting mapping $\tau$ is $1$-, 22- and $E$-coherent and $1$-, 22- and $E$-saturated on $W$.
@@
For each $E^{0}$-equivalence class $T$ with $h(T)\geq 0$, consider the $\Theta^{1}(\Phi)$-quasistate $\boldsymbol{q}^{T}\in\boldsymbol{Q}_{1}(\varphi)$ such that $\mathfrak{M}^{0},v\models(\hat{\boldsymbol{q}}^{T})^{\mathbf{1}}$ for all (equivalently, some) $v\in T$. Observe that $\boldsymbol{q}^{T}$ may contain formulas $\chi$ with $\textit{md}^{\mathbf{1}}{(\chi)}>h(T)$. Let $\Phi^{T}=\Theta^{\mathbf{1}}(\Phi)\cap\Gamma^{\mathbf{1}}_{h(T)}(\Phi)$ and consider the restriction $\boldsymbol{q}_{0}^{T}=\{t|_{\smash{\Gamma^{\mathbf{1}}_{h(T)}(\Phi)}}\mid t\in\boldsymbol{q}^{T}\}$, which is a $\Phi^{T}$-quasistate. Denote $\smash{\hat{\boldsymbol{q}}^{T}_{0}}$ by $\varphi^{T}$. Since $\textit{adp}^{\mathbf{2}}(\varphi^{T})<k$, we apply our induction hypothesis to the $\mathcal{L}_{1,2,E}$-formula $\varphi^{T}$ and obtain a tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}=((W^{T},R_{1}^{T},R_{2}^{T},E^{T}),T,\tau^{T})$ for $\mathop{\textit{sub}}\varphi^{T}$ (equivalently, for $\Phi^{T}$) such that $\tau^{0}(v)|_{\Theta^{\mathbf{1}}(\Phi)}=\tau^{T}(v)$ for all $v\in T$. At each $E^{0}$-equivalence class $T$ with $h(T)<0$, we take a degenerate tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}$ for $\emptyset$. Next, for each $T\neq T_{0}$, we graft the tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}$ onto $T$ in the $(1,\kappa)$-quasimodel $\mathfrak{Q}$ to obtain a tapered $(1,\kappa)$-cactus $\mathfrak{C}$. A fully grafted tapered $(1,\kappa)$-cactus is obtained by grafting, in addition, $\mathfrak{C}^{T^{0}}$ onto $T^{0}$ in $\mathfrak{C}$.